# Targeted Deep Learning System Boundary Testing

### v2

@@ 3.3. Optimizer
We combine the latent vector of the initial class $w_{0}$ and the latent vector of a target class $w_{g}$ into a new latent vector $w^{\prime}=\kappa w_{0}+(1-\kappa)w_{g}$ (see 4(b)), where $\kappa$ is the manipulation strategy found by the optimizer. To find adequate manipulation strategies, Mimicry uses the AGE-MOEA-2 optimizer (Panichella, 2022), known for its outstanding performance with one or multiple objectives. Contrary to the original style mixing approach of Karras et al. (Karras et al., 2019b), the produced strategies $\kappa$ do not swap individual layers (4(a)), but they are weights for linear interpolation between two layers of the same rank in two latent vectors (4(b)). This increases the controllability of feature mixing between the source and target seed, which is useful for precise DL boundary assessment.
@@ 3.4. Objectives for Boundary Testing
In our experiments, we use two objectives to optimize toward boundary candidates. The first objective function consists in a dynamic confidence balance, depicted in (missing) 5. Here $\hat{y}^{\prime}$ are the predicted class probabilities of the manipulated input $X^{\prime}$. The subscript $1st$ denotes the index of the primary seed class and the set $J=\mathcal{C}_{g}$ are all the other elements that should be considered (in our case, we only consider the second most-likely class from the SUT). This function essentially quantifies how similar multiple confidences are to each other and how much weight they have combined against the rest of the confidence values. This means that the more similar the targeted confidence values are and the more confidence they encapsulate in the prediction, the higher the dynamic confidence balance.
(5) $\omega_{dcb}:\frac{\sum_{j\in J}|\hat{y}^{\prime}_{1}-\hat{y}^{\prime}_{j}|}{||J||(\hat{y}^{\prime}_{1}+\sum_{j\in J}\hat{y}^{\prime}_{j})}.$
In addition to (missing) 5, we also integrate a quality criterion, similar to DeepJanus (Riccio and Tonella, 2020). This quality criterion measures the normalized Euclidean distance $d_{2}$ between strategies (genomes) of an archive $\kappa_{A}$ and a population $\kappa_{P}$, to enforce a greater novelty of solutions in genomes. Here we want to maximize sparsity of new genomes, based on the parent population which functions as the archive, therefore we minimize (missing) 6. The search for novelty is commonly used to limit the exploitation of local minima and have a better traversal of the optimizer search space. This novelty measure does not concern the actual generated images, rather their manipulation weights.
(6) $\omega_{d2}:1-min\Big{\{}\frac{d_{2}(a,p)}{\sqrt{||a||}}\;\Big{|}\;(a,p)\in\kappa_{A}\times\kappa_{P}\Big{\}}.$
@@ 4.2.1. Datasets
In our study, we used five image classification datasets, namely MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), SVHN (Netzer et al., 2011), CIFAR-10 (Krizhevsky et al., 2009), and ImageNet-1k (Deng et al., 2009) (hereafter referred to as ImageNet for simplicity of exposition). We chose these five datasets because three (MNIST, FashionMNIST and SVHN) are compatible with the both our baselines DeepJanus (Riccio and Tonella, 2020, 2023) and Sinvad (Kang et al., 2024, 2020). This selection is also consistent with previous studies, such as Dola et al. (Dola et al., 2024). However, Mimicry can be applied to any image dataset. CIFAR-10 and ImageNet are used to demonstrate the generalizability of our approach to data sets where a model input representation is not available for DeepJanus and, thus, we compare against the Sinvad generative-based approach (Kang et al., 2024) as a baseline.
MNIST. Dataset of handwritten digits (LeCun et al., 1998) consisting of grayscale images $28\times 28$ labeled with the corresponding digit (the possible classes range from 0 to 9). MNIST has 60,000 training inputs and 10,000 test inputs. As StyleGAN only allows square images of size $2^{n}$, we zero pad the images to scale them to size and duplicate channels resulting in an image of size $32\times 32\times 3$.
FashionMNIST. Another dataset consisting of $28\times 28$ grayscale images of Zalando’s articles belonging to $10$ categories (Xiao et al., 2017). The dataset has more complex patterns and variations than MNIST and contains $60,000$ images for training and $10,000$ for testing. Similarly to MNIST we again zero pad the data to make it compatible with StyleGAN having a shape of $32\times 32\times 3$.
SVHN. A more complex dataset contains $32\times 32\times 3$ color digits of house numbers cropped from Google Street View images (Netzer et al., 2011). As the data is already compatible with StyleGANs no transformation were used. It has $73,257$ training input and $26,032$ test input. The classification task is particularly challenging due to variations in lighting, background clutter, and the presence of distracting digits adjacent to the digit of interest.
CIFAR-10. Another standard benchmark for image classification tasks is divided into $10$ classes of different objects (Krizhevsky et al., 2009) and divided into $50,000$ training images and $10,000$ testing images. Although the images are small ($32\times 32\times 3$), they contain visual complexities and variations of real-world objects, requiring models to extract meaningful features from low-resolution images. Again the data is compatible with StyleGAN as such no transformations were performed.
@@ 4.2.2. System Under Test
We adopt a WideResNet-50-2 classifier (Zagoruyko, 2016) available in the PyTorch library (Paszke et al., 2017) with pre-trained weights for ImageNet. The model achieves an accuracy of $0.816$ on ImageNet1k. For the other datasets (MNIST, FashionMNIST, SVHN, CIFAR-10), we train the same WideResNet-50-2 architecture with the default train data splits given in Torchvision (maintainers and contributors, 2016). We prioritized consistency across SUTs by using the same model architecture for all datasets, rather than using state-of-the-art or literature-sourced models, in order to reduce variability in our results due to architectural peculiarities. The trained networks achieved an accuracy above 0.9 on the test split for all datasets, except for CIFAR-10, where the accuracy was 0.81. For optimization in training, we use AdamW, which has demonstrated superior performance across multiple tasks due to its adaptive weight decay compared to traditional Adam (Loshchilov and Hutter, 2017). Furthermore, we schedule our learning rates using the OneCycle strategy (Smith and Topin, 2018), which initially increases the learning rate toward a maximum value and then decreases it again to a set minimum, forming a learning rate trajectory similar to a right-skewed Gaussian. This type of scheduling has been shown to improve convergence during training (Smith and Topin, 2018).
@@ 4.3. Metrics
Boundary Distance ($\downarrow$). In order to quantify whether a candidate is a good boundary input, we measure the Euclidean distance $d_{2}$ between the predicted classes $\hat{y}^{\prime}\in\mathbb{R}^{||\mathcal{C}||}$ to the theoretical boundary (missing) 7. In this work, we specifically look at identifying boundary candidates between two classes rather than multiple, the theoretical boundary is assumed to be a vector $b\in\mathbb{R}^{||\mathcal{C}||}$ where $\sum b=1$, and all non-zero elements are equal to $\frac{1}{1+||C_{g}||}$, that is in equilibrium between the classes used for boundary discovery $C_{g}$ and the original class $C_{0}$.The lower this measure, the better the boundary input.
@@
Label Coverage ($\uparrow$) & Escape Ratio ($\downarrow$). Another important aspect of test case generation is the coverage of possible test outcomes. The Label Coverage indicates the distribution of the target labels, $\mathcal{Y}^{\prime}_{t}=\{\hat{y}^{\prime}_{t}\forall X^{\prime}\}$, for the candidates of the boundaries, measured using the Kolmogorov–Smirnov distance (missing) 8. Here, $\mathcal{U}_{t}=\mathcal{U}\{\mathcal{C}\;\textbackslash\;\{c_{0}\}\}$ represents the uniform distribution of all possible target labels. A value of 1 indicates a uniform distribution in all possible classes, while a value of 0 indicates that all test cases share the same target label. Although achieving a perfect value of 1 is often infeasible due to the spatial separation of some classes on the decision surface, higher label coverage is generally preferred.
@@
(9) $m_{4}(\mathcal{P},\mathcal{Y},\mathcal{Y}^{\prime})=\frac{1}{||\mathcal{P}||}\sum_{\hat{y}\in\mathcal{Y},\hat{y}^{\prime}\in\mathcal{Y}^{\prime}}[\hat{y}_{1st}\notin\{\hat{y}^{\prime}_{1st},\hat{y}^{\prime}_{2nd}\}].$
Laplacian Variance of Image Differences ($\uparrow$). This measure aims to quantify the change in information between the boundary input and the initial input (missing) 10. Essentially, it indicates whether the method produces functionally different outputs or merely corrupts or blurs an image, common phenomena observed when applying simple pertubations to the latent space. Here, $L$ is a $3\times 3$ Laplacian kernel, applied using convolution on the differences of two images. The higher this measure, the better the boundary input.
@@
Concerning RQ3, we performed an evaluation study with human assessors to evaluate the quality of the generated inputs. Quality was assessed with several characteristics, described as follows. Label preservation describes whether the original class label is still assigned to the generated candidate by the evaluator. The inverse of which is target preservation, which quantifies if the boundary target is visually depicted in the generated image. When combining these two we get boundary preservation, as the generated image shows visual elements of either classes of the boundary. The latter is simply the sum of the first two. Finally, we measure the validity (Riccio and Tonella, 2023), i.e., whether any class within the considered domain was associated by the evaluator, meaning the image still has valid syntactic features to the human observer.
@@ 4.4. Baseline Approaches
Sinvad (Kang et al., 2024) is a latent manipulation-based tool that leverages Variational Autoencoders (VAEs) as its generative networks. Specifically, Sinvad encodes test inputs into latent vectors using a VAE trained on the corresponding training set. Since the approach relies on population-based optimization, the initial population is derived from the latent vector of the original images, perturbed with random noise to simulate a diverse set of slightly altered inputs. Sinvad then optimizes toward a fitness function through uniform crossover of latent vectors and mutation via noise. The mutation severity is dynamically adjusted: if progress toward the objective function stalls, the mutation magnitude decreases accordingly. In our study we compare Sinvad against Mimicry on all datasets using the pre-trained VAEs available in the replication package (Kang et al., 2024).
DeepJanus (Riccio and Tonella, 2020) is a representative of model-based approaches and uses a multi-objective search-based algorithm to mutate the control points of a model of the inputs, to generate pairs of inputs that are close to each other, yet produce different behaviors of the DL system (Riccio and Tonella, 2020). The input model representation is obtained through a vectorization operation, which produces a sequence of control points that are iteratively displaced to achieve slight modifications. The input image can then be reconstructed through a rasterization operation. Our comparison focused on the MNIST, FashionMNIST, and SVHN datasets, as DeepJanus’s model representation supports these benchmarks. For CIFAR-10 and ImageNet, DeepJanus is not applicable since an appropriate input model is not available for such a feature-rich dataset, and cannot be created with the adopted vectorization-rasterization approach, as noted by its authors (Riccio and Tonella, 2020, 2023).
@@ 4.5. Procedure
Our approach requires generating seeds by sampling latent vectors using a conditional StyleGAN architecture. For MNIST, FashionMNIST, and SVHN, pre-trained conditional StyleGAN2 networks are not available. Thus, we trained them on the training partition of each dataset, following the training configurations and guidelines of the original paper (Karras et al., 2019a). To monitor the model’s performance during training, we used the Fréchet Inception Distance (FID) metric (Heusel et al., 2017). The final FID score obtained is 0.91 for MNIST, 2.34 for FashionMNIST and 4.20 for SVHN, which is in line with the original paper (Karras et al., 2019a). For CIFAR-10, we used pre-trained StyleGAN2 networks available in the literature (Karras et al., 2020a). For ImageNet, we used a pre-trained StyleGAN-XL (Sauer et al., 2022).
For RQ1, for all applicable datasets, we execute all test generators (Mimicry, Sinvad, DeepJanus) using a budget of $15,000$ predictions of the SUT per boundary candidate to be optimized for. Given the varying approaches of the methods, we define the SUT as the determinant of the budget, ensuring consistency across experiments. While the budget may be reached, it does not need to be fully utilized, as some methods may terminate earlier. Particularly, we instruct the tools to search for 10 boundary candidates for each class, giving us a total budget of $1.5M$ predictions per test method and dataset. Overall, our study includes nearly $20M$ predictions ($2$ tools, Mimicry and Sinvad $\times$ $1.5M$ predictions $\times$ $5$ datasets + $1.5M$$\times$ $3$ datasets for DeepJanus).
@@
For RQ3, as image metrics often do not coincide with human perception (Lambertenghi and Stocco, 2024), we use a human evaluation study to quantify the quality of generated cases. The evaluators are recruited on AWS Mechanical Turk, with an attention question incorporated to filter out inattentive or disengaged responses (Sorokin and Forsyth, 2008). The attention question was given with the datasets reference images, asking the participants to select a specific class. Prior to the evaluation, images of the original datasets were shown to make the assessors familiar with the classes. For each dataset, we randomly selected 10 generated images from each method (Mimicry, Sinvad, DeepJanus), covering all classes. We ask the participants to select all classes they can recognize in the provided image, with an “Not applicable” option if they could not identify any of the provided classes. Each participant was only shown samples from a single dataset. We recruited 30 distinct evaluators for each dataset, with $\sim 20$ valid respondents per dataset, as some had to be discarded due to incorrect attention question answers.
@@ 4.6.1. RQ1 (effectiveness)
Boundary distance 0.283 1.19e-7$\bullet$ 9.51e-7$\bullet$ 0.005$\bullet$ 5.21e-9 $\bullet$ 7.7e-15 $\bullet$ 4.08e-31 $\bullet$ $-1$bullet
@@
Label coverage 0.986 0.998 0.998 0.32 0.213 0.006 $\bullet$ 0.001$\bullet$ 1.41e-40 $\bullet$
Aggregating these measures, we are interested in whether these differences have statistical significance. Therefore we employ a one-tailed Mann–Whitney U test (Wilcoxon, 1945) (with $\alpha=0.05$) between Mimicry and the baseline methods. Additionally, we calculate the Cohen’s $d$ effect size (Cohen, 1988), whose magnitude is indicated in Table 2 by a colored bullet ($\bullet$), where ${\color[rgb]{0,0.52,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.52,0}\bullet}=d>1$ (large) $,\;{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bullet}=1\geq d>0.5$ (medium) $,\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}=d\leq 0.5$ (small). The statistical results confirm the trend observed in the figures, where Mimicry performs well on all datasets and outperforms the baselines especially as their complexity increases.

### v3

@@ 3.3. Optimizer
We combine the latent vector of the initial class $w_{0}$ and the latent vector of a target class $w_{g}$ into a new latent vector $w^{\prime}=\kappa w_{0}+(1-\kappa)w_{g}$ (see 4(b)), where $\kappa$ is the manipulation strategy found by the optimizer. To find adequate manipulation strategies, Mimicry uses the AGE-MOEA-2 optimizer (panichella2022improved), known for its outstanding performance with one or multiple objectives. Contrary to the original style mixing approach of Karras et al. (karras2019style), the produced strategies $\kappa$ do not swap individual layers (4(a)), but they are weights for linear interpolation between two layers of the same rank in two latent vectors (4(b)). This increases the controllability of feature mixing between the source and target seed, which is useful for precise DL boundary assessment.
@@ 3.4. Objectives for Boundary Testing
In our experiments, we use two objectives to optimize toward boundary candidates. The first objective function consists in a dynamic confidence balance, depicted in () 5. Here $\hat{y}^{\prime}$ are the predicted class probabilities of the manipulated input $X^{\prime}$. The subscript $1st$ denotes the index of the primary seed class and the set $J=\mathcal{C}_{g}$ are all the other elements that should be considered (in our case, we only consider the second most-likely class from the SUT). This function essentially quantifies how similar multiple confidences are to each other and how much weight they have combined against the rest of the confidence values. This means that the more similar the targeted confidence values are and the more confidence they encapsulate in the prediction, the higher the dynamic confidence balance.
[…]
In addition to () 5, we also integrate a quality criterion, similar to DeepJanus (riccio2020deepjanus). This quality criterion measures the normalized Euclidean distance $d_{2}$ between strategies (genomes) of an archive $\kappa_{A}$ and a population $\kappa_{P}$, to enforce a greater novelty of solutions in genomes. Here we want to maximize sparsity of new genomes, based on the parent population which functions as the archive, therefore we minimize () 6. The search for novelty is commonly used to limit the exploitation of local minima and have a better traversal of the optimizer search space. This novelty measure does not concern the actual generated images, rather their manipulation weights.
[…]
@@ 4.2.1. Datasets
In our study, we used five image classification datasets, namely MNIST (mnist), FashionMNIST (fashionmnist), SVHN (svhn), CIFAR-10 (cifar), and ImageNet-1k (deng2009imagenet) (hereafter referred to as ImageNet for simplicity of exposition). We chose these five datasets because three (MNIST, FashionMNIST and SVHN) are compatible with the both our baselines DeepJanus (riccio2020deepjanus; 2023-Riccio-ICSE) and Sinvad (sinvad-tosem; kang2020sinvad). This selection is also consistent with previous studies, such as Dola et al. (dola2024cit4dnn). However, Mimicry can be applied to any image dataset. CIFAR-10 and ImageNet are used to demonstrate the generalizability of our approach to data sets where a model input representation is not available for DeepJanus and, thus, we compare against the Sinvad generative-based approach (sinvad-tosem) as a baseline.
MNIST. Dataset of handwritten digits (mnist) consisting of grayscale images $28\times 28$ labeled with the corresponding digit (the possible classes range from 0 to 9). MNIST has 60,000 training inputs and 10,000 test inputs. As StyleGAN only allows square images of size $2^{n}$, we zero pad the images to scale them to size and duplicate channels resulting in an image of size $32\times 32\times 3$.
FashionMNIST. Another dataset consisting of $28\times 28$ grayscale images of Zalando’s articles belonging to $10$ categories (fashionmnist). The dataset has more complex patterns and variations than MNIST and contains $60,000$ images for training and $10,000$ for testing. Similarly to MNIST we again zero pad the data to make it compatible with StyleGAN having a shape of $32\times 32\times 3$.
SVHN. A more complex dataset contains $32\times 32\times 3$ color digits of house numbers cropped from Google Street View images (svhn). As the data is already compatible with StyleGANs no transformation were used. It has $73,257$ training input and $26,032$ test input. The classification task is particularly challenging due to variations in lighting, background clutter, and the presence of distracting digits adjacent to the digit of interest.
CIFAR-10. Another standard benchmark for image classification tasks is divided into $10$ classes of different objects (cifar) and divided into $50,000$ training images and $10,000$ testing images. Although the images are small ($32\times 32\times 3$), they contain visual complexities and variations of real-world objects, requiring models to extract meaningful features from low-resolution images. Again the data is compatible with StyleGAN as such no transformations were performed.
@@ 4.2.2. System Under Test
We adopt a WideResNet-50-2 classifier (zagoruyko2016wide) available in the PyTorch library (paszke2017automatic) with pre-trained weights for ImageNet. The model achieves an accuracy of $0.816$ on ImageNet1k. For the other datasets (MNIST, FashionMNIST, SVHN, CIFAR-10), we train the same WideResNet-50-2 architecture with the default train data splits given in Torchvision (torchvision2016). We prioritized consistency across SUTs by using the same model architecture for all datasets, rather than using state-of-the-art or literature-sourced models, in order to reduce variability in our results due to architectural peculiarities. The trained networks achieved an accuracy above 0.9 on the test split for all datasets, except for CIFAR-10, where the accuracy was 0.81. For optimization in training, we use AdamW, which has demonstrated superior performance across multiple tasks due to its adaptive weight decay compared to traditional Adam (loshchilov2017decoupled). Furthermore, we schedule our learning rates using the OneCycle strategy (smith2018superconvergence), which initially increases the learning rate toward a maximum value and then decreases it again to a set minimum, forming a learning rate trajectory similar to a right-skewed Gaussian. This type of scheduling has been shown to improve convergence during training (smith2018superconvergence).
@@ 4.3. Metrics
[…]
@@
[…]
@@
In addition to label coverage, the escape ratio quantifies the fraction of test cases that no longer consider the initial class as the origin of the boundary () 9. In this case, $\mathcal{Y}$ represents the set of predictions on all initial seeds, $\mathcal{Y}^{\prime}$ is the set of predictions for the candidates generated, and $\left[\cdot\right]$ is the Iverson bracket. The subscript on the predictions denotes the indices taken when an argmax is applied to the vector. This measure is critical because, when testing boundaries for a specific class, we are interested only in boundaries that relate to the original class. Thus, the escape ratio should be small to ensure that the generated candidates remain relevant to the objective.
@@
Laplacian Variance of Image Differences ($\uparrow$). This measure aims to quantify the change in information between the boundary input and the initial input () 10. Essentially, it indicates whether the method produces functionally different outputs or merely corrupts or blurs an image, common phenomena observed when applying simple pertubations to the latent space. Here, $L$ is a $3\times 3$ Laplacian kernel, applied using convolution on the differences of two images. The higher this measure, the better the boundary input.
@@
Concerning RQ3, we performed an evaluation study with human assessors to evaluate the quality of the generated inputs. Quality was assessed with several characteristics, described as follows. Label preservation describes whether the original class label is still assigned to the generated candidate by the evaluator. The inverse of which is target preservation, which quantifies if the boundary target is visually depicted in the generated image. When combining these two we get boundary preservation, as the generated image shows visual elements of either classes of the boundary. The latter is simply the sum of the first two. Finally, we measure the validity (2023-Riccio-ICSE), i.e., whether any class within the considered domain was associated by the evaluator, meaning the image still has valid syntactic features to the human observer.
@@ 4.4. Baseline Approaches
Sinvad (sinvad-tosem) is a latent manipulation-based tool that leverages Variational Autoencoders (VAEs) as its generative networks. Specifically, Sinvad encodes test inputs into latent vectors using a VAE trained on the corresponding training set. Since the approach relies on population-based optimization, the initial population is derived from the latent vector of the original images, perturbed with random noise to simulate a diverse set of slightly altered inputs. Sinvad then optimizes toward a fitness function through uniform crossover of latent vectors and mutation via noise. The mutation severity is dynamically adjusted: if progress toward the objective function stalls, the mutation magnitude decreases accordingly. In our study we compare Sinvad against Mimicry on all datasets using the pre-trained VAEs available in the replication package (sinvad-tosem).
DeepJanus (riccio2020deepjanus) is a representative of model-based approaches and uses a multi-objective search-based algorithm to mutate the control points of a model of the inputs, to generate pairs of inputs that are close to each other, yet produce different behaviors of the DL system (riccio2020deepjanus). The input model representation is obtained through a vectorization operation, which produces a sequence of control points that are iteratively displaced to achieve slight modifications. The input image can then be reconstructed through a rasterization operation. Our comparison focused on the MNIST, FashionMNIST, and SVHN datasets, as DeepJanus’s model representation supports these benchmarks. For CIFAR-10 and ImageNet, DeepJanus is not applicable since an appropriate input model is not available for such a feature-rich dataset, and cannot be created with the adopted vectorization-rasterization approach, as noted by its authors (riccio2020deepjanus; 2023-Riccio-ICSE).
@@ 4.5. Procedure
Our approach requires generating seeds by sampling latent vectors using a conditional StyleGAN architecture. For MNIST, FashionMNIST, and SVHN, pre-trained conditional StyleGAN2 networks are not available. Thus, we trained them on the training partition of each dataset, following the training configurations and guidelines of the original paper (DBLP:journals/corr/abs-1912-04958). To monitor the model’s performance during training, we used the Fréchet Inception Distance (FID) metric (DBLP:journals/corr/HeuselRUNKH17). The final FID score obtained is 0.91 for MNIST, 2.34 for FashionMNIST and 4.20 for SVHN, which is in line with the original paper (DBLP:journals/corr/abs-1912-04958). For CIFAR-10, we used pre-trained StyleGAN2 networks available in the literature (DBLP:journals/corr/abs-2006-06676). For ImageNet, we used a pre-trained StyleGAN-XL (sauer2022styleganXL).
For RQ1, for all applicable datasets, we execute all test generators (Mimicry, Sinvad, DeepJanus) using a budget of $15,000$ predictions of the SUT per boundary candidate to be optimized for. Given the varying approaches of the methods, we define the SUT as the determinant of the budget, ensuring consistency across experiments. While the budget may be reached, it does not need to be fully utilized, as some methods may terminate earlier. Particularly, we instruct the tools to search for 10 boundary candidates for each class, giving us a total budget of $1.5M$ predictions per test method and dataset. Overall, our study includes nearly $20M$ predictions ($2$ tools, Mimicry and Sinvad $\times$ $1.5M$ predictions $\times$ 55 datasets + $1.5M$ $\times$ 33 datasets for DeepJanus).
@@
For RQ3, as image metrics often do not coincide with human perception (lambertenghi2024assessing), we use a human evaluation study to quantify the quality of generated cases. The evaluators are recruited on AWS Mechanical Turk, with an attention question incorporated to filter out inattentive or disengaged responses (sorokin2008utility). The attention question was given with the datasets reference images, asking the participants to select a specific class. Prior to the evaluation, images of the original datasets were shown to make the assessors familiar with the classes. For each dataset, we randomly selected 10 generated images from each method (Mimicry, Sinvad, DeepJanus), covering all classes. We ask the participants to select all classes they can recognize in the provided image, with an “Not applicable” option if they could not identify any of the provided classes. Each participant was only shown samples from a single dataset. We recruited 30 distinct evaluators for each dataset, with $\sim 20$ valid respondents per dataset, as some had to be discarded due to incorrect attention question answers.
@@ 4.6.1. RQ1 (effectiveness)
[…]
@@
[…]
Aggregating these measures, we are interested in whether these differences have statistical significance. Therefore we employ a one-tailed Mann–Whitney U test (Wilcoxon1945) (with $\alpha=0.05$) between Mimicry and the baseline methods. Additionally, we calculate the Cohen’s $d$ effect size (cohen1988statistical), whose magnitude is indicated in Table 2 by a colored bullet ($\bullet$), where ${\color[rgb]{0,0.52,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.52,0}\bullet}=d>1$ (large) ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bullet}=1\geq d>0.5$ (medium) ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}=d\leq 0.5$ (small). The statistical results confirm the trend observed in the figures, where Mimicry performs well on all datasets and outperforms the baselines especially as their complexity increases.